OpenStudy (anonymous):

Kevin is trying to find a brown sock in his drawer. He has 16 white socks, 4 brown socks, and 6 black socks. What is the probability that he pull out either a black or white sock, puts it back and then pulls out a brown sock?

OpenStudy (anonymous):

there are 26 socks in total (16 + 4 + 6) so we are asked if if he pulls out a black OR white sock. the or should signal to you that you add these probabilities \[16/26 + 6/26 = 22/26\] now, it says that he puts it back in (in math terms, he replaces it) and then picks out a brown sock \[4/26\] now we just multiply those two probabilities together. the key word you want to look for is AND. and implies that you multiply the probabilites \[22/26*4/26 = 22/13*1/13 = 22/169\]

OpenStudy (anonymous):

Two six-sided dice are rolled and the their SUM is recorded. The number of possible outcomes for this event is